Simplify. (a + b) (c – d) + (a – b) (c + d) + 2 (ac + bd)

Simplify. 
 (a + b) (c – d) + (a – b) (c + d) + 2 (ac + bd)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a + b) (c – d) + (a – b) (c + d) + 2 (ac + bd)
= (a + b) (c–d) + (a–b) (c + d) + 2 (ac + bd)
=(ac - ad + bc - bd) + (ac + ad - bc - bd) + (2ac + 2bd)
= ac - ad + bc - bd + ac + ad - bc - bd + 2ac + 2bd= 4ac

Thanks